OpenStudy (anonymous):

4(7x - 1) = 3(2x + 1) - 4(x + 6)

OpenStudy (anonymous):

28x-4=6x+3-4x-24 28x-6x+4x=3-24+4 26x=-17 x=-17/26

OpenStudy (saifoo.khan):

-17/16

OpenStudy (aroub):

28x-4=6x+3-4x+24 x=-17/26 , -17/16